Water flowing over the Wraith Falls in Yellowstone National Park drops 30 m.

 If all of the water's energy goes to heat it, what is its temperature increase?


1
Expert's answer
2022-03-03T13:23:09-0500

Let all potential energy be converted to heat energy:


"mgh=mc\Delta T"

where "m" is the mass of water, "g = 9.8m\/s^2" is the gravitational acceleration, "h=30m" is the height, "c = 4200J/kg/\degree C" is the specific heat capacity of water, "\Delta T" increase in temperature. Thus, obtain:


"\Delta T = \dfrac{gh}{c} = \dfrac{9.8m/s^2\cdot 30m}{4200J/kg/\degree C} = 0.07\degree C"

Answer. "0.07\degree C".